OpenStudy (anonymous):

use the graph of the integrand and areas to evaluate the integral: upper limit 1, lower limit -1, f(x)=(2-lxl). can you explain how to evaluate this intergral?

OpenStudy (mathteacher1729):

I guess step 1 is drawing a graph ?

OpenStudy (anonymous):

I have the graph on calculator, but I am not sure how to determine area without using individual rectangles

OpenStudy (mathteacher1729):

You should be able to do this using triangles.

OpenStudy (slaaibak):

Use the fact that it's symmetric. Only evaluate from 1 to 0, and multiply by 2

OpenStudy (anonymous):

thanks let me go work with it and see if i can get the answer that the book has

OpenStudy (anonymous):

f(x)=2-(-x )=2+x, when -1<x<0 similarly f(x)=2-x , when 0<=x<1 so integrate these two and add them , you will get the ans

OpenStudy (anonymous):

ans is 3

OpenStudy (anonymous):

I got it but i ended up having too just divide it into two intervals and solve the area of these two intervals

OpenStudy (anonymous):

yes you got it